suppose you borrow $1500 at 12% simple interest for 30 months. How much money would you owe at the end of that time ?
Answer by stanbon(75887)

suppose you borrow $1500 at 12% simple interest for 30 months. How much money would you owe at the end of that time ?
-----------------------
A = P + Prt = P(1 + rt)
-----------------
r = 0.12
t = 30/12 = 2.5
-------------------
A = 1500(1 + 0.12*2.5)
A = 1500*1.3
A = $1950
===============
Cheers,
Stan H.